One liter per minute of liquid containing A and B (CAo=0.10 mole/liter, CBo= 0.01mole/liter) flow into a mixed flow reactor of volume V-1 liter. The materials react in a complex manner for which the stoichiometry is unknown. The outlet stream from the reactor contains A, B and C (CA 0.02 mole liter, Cu 0.03 mole liter, Cer=0.04 mole/liter). Find the rate of reaction of A, B and C (ie, rA,rB,rC and to) for the conditions within the reactor


1
Expert's answer
2021-09-02T02:33:00-0400

"k_A = \dfrac{c_{iA}}{c_{oA}} = \dfrac{0.10}{0.02} = 5"


"k_B = \dfrac{c_{iB}}{c_{B}} = \dfrac{0.01}{0.03} = 1/3"


"k_C = \dfrac{c_{iC}}{c_{oC}} = \dfrac{0.01}{0.06} =1/ 6"